### v2

# Robustness Analysis of USmorph: I. Generalization Efficiency of Unsupervised Strategies and Supervised Learning in Galaxy Morphological Classification

Abstract
We conduct a systematic robustness analysis of the hybrid machine learning framework USmorph, which integrates unsupervised and supervised learning for galaxy morphological classification. Although USmorph has already been applied to nearly 100,000 $I$-band galaxy images in the COSMOS field ($0.2<z<1.2$, $I_{\mathrm{mag}}<25$), the stability of its core modules has not been quantitatively assessed. Our tests show that the convolutional autoencoder (CAE) achieves the best performance in preserving structural information when adopting an intermediate network depth, $5\times 5$ convolutional kernels, and a 40-dimensional latent representation. The adaptive polar coordinate transform (APCT) effectively enhances rotational invariance and improves the robustness of downstream tasks. In the unsupervised stage, a bagging clustering number of $K=50$ provides the optimal trade-off between classification granularity and labeling efficiency. For supervised learning, we employ GoogLeNet, which exhibits stable performance without overfitting. We validate the reliability of the final classifications through two independent tests: (1) the t-distributed stochastic neighbor embedding (t-SNE) visualization reveals clear clustering boundaries in the low-dimensional space; and (2) the morphological classifications are consistent with theoretical expectations of galaxy evolution, with both true and false positives showing unbiased distributions in the parameter space. These results demonstrate the strong robustness of the USmorph algorithm, providing guidance for its future application to the China Space Station Telescope (CSST) mission.
Galaxy structure (622), Astrostatistics techniques (1886), Astronomy data analysis (1858)
I Introduction
The morphological characteristics of galaxies are crucial to understanding their formation and evolution. Galaxies exhibit a rich diversity in morphology, including features such as spiral arms, bars, rings, and tidal tails (e.g.,Ho+2019; Sofue+2021; ren+2023; Wei+2024). These large-scale structures provide critical insights into galaxies’ interactions, ongoing physical processes, and evolutionary pathways, as reflected by their physical properties like colors, redshifts, masses, and environmental factors (e.g., Kauffmann+2004; omand+2014; schawinski+2014; kawinwanichakij+2017; Gu+2018; su+2025). Consequently, they offer important clues for a comprehensive understanding of galaxy populations and their evolution.
In early citizen science projects such as Galaxy Zoo (GZ; Lintott+2008; Lintott+2021), galaxies were classified by visual inspection. With the development of technology, both parametric and non-parametric methods have been introduced for galaxy classification. Parametric modeling employs predefined mathematical functions such as the Sérsic profile (S+1963; Balcells+2003) to parameterize the two-dimensional surface brightness distribution of galaxies, allowing the extraction of structural parameters like the effective radius ($r_{\mathrm{e}}$), Sérsic index ($n$), and the axis ratio ($b/a$). Parametric methods typically assume that galaxies possess a single, symmetric, and smooth light distribution. In contrast, non-parametric methods avoid assumptions about functional forms and instead compute statistical indicators directly from image pixels to characterize the light distribution. Examples include the CAS system (Concentration $C$, Asymmetry $A$, and Smoothness $S$; Conselice+2000; Conselice+2003), Gini coefficient, $M_{20}$, and $MID$ statistics (e.g., Lotz+2004; Lotz+2006; Freeman+2013). These indicators quantify the complexity and irregularity of galaxy morphologies, providing objective measures of their structure.
Visual inspection, parametric, and non-parametric methods have proven effective in past large-scale sky surveys. However, visual inspection is time-consuming and labor-intensive, parametric fitting relies on idealized assumptions, and non-parametric statistics have limited descriptive power. These drawbacks limit their application to large-scale sky surveys in the new era. Recent development of large-scale sky surveys, including the Sloan Digital Sky Survey (SDSS; Stoughton+2002), the Legacy Survey of Space and Time (LSST; Abell+2009), the Euclid space telescope (Euclid; Euclid+2025), and the forthcoming China Space Station Telescope (CSST; CSST+2025), will generate vast amounts of high-resolution, multi-band galaxy images. The implementation of machine learning (ML) algorithms for automated galaxy image classification therefore becomes paramount.
Convolutional neural networks (CNNs; sun+2018) have shown remarkable ability in the automated classification of galaxy morphologies, as they can effectively capture structural and textural details in images, enabling accurate and efficient identification (e.g. Krizhevsky+2012; Davis+2014; Dieleman+2015; Fernando+2024). However, as a supervised learning method (SML), the performance of CNNs heavily relies on both the quality and the size of the labeled training dataset. Currently, data annotation remains laborious and costly, mainly dependent on manual labeling. For example, the well-known ImageNet project (Deng+2009) took nearly a decade to complete the annotation of approximately 12 million images. Therefore, unsupervised machine learning (UML) has been widely adopted for the analysis of large-scale unlabeled galaxy datasets since it does not rely on labeled data. UML techniques can directly extract features from galaxy images while simultaneously uncovering intrinsic data structures and patterns, allowing for clustering based on morphological similarity ((Hocking+2018; Kolesnikov+2023; Tohill+2024)). Recently, Kolesnikov+2023 applied the SOMBRERO algorithm, an unsupervised clustering method based on EGG morphological features, to classify galaxies in both SDSS and HST datasets.
However, the performance of UML methods heavily depends on the quality of the feature representations. In high-dimensional data domains such as images, feature structures are usually highly complex and typically contain significant noise and redundant information (e.g., xue+2023). Moreover, many UML methods rely on a single clustering strategy, which may yield inconsistent results when different similarity metrics, such as centroid-based, hierarchical, or density-based measures, are employed. These inconsistencies can lead to classification errors and undermine the reliability of the clustering process.
To overcome these limitations, Song+2024 proposed the USmorph hybrid machine-learning framework, which integrates feature extraction, unsupervised clustering, and supervised classification to reduce reliance on labeled data. The workflow involves: (1) employing a convolutional autoencoder (CAE; Massey+2009) to denoise and compress galaxy images; (2) applying an adaptive polar coordinate transformation (APCT; Fang+2023) to enhance the model’s rotational invariance; and (3) using a bagging-based voting clustering method (Zhou+2022) to group features and generate training labels for a supervised CNN classifier (GoogLeNet; szegedy+2015). The supervised learning phase employs the GoogLeNet model, which achieved an overall accuracy of approximately 94% on the validation set. The classification outcomes align closely with the parametric morphological measurements and non-parametric morphological spatial distributions of galaxies, thereby confirming the classification’s reliability. With USmorph, Song+2024 successfully classified nearly $\sim$100,000 galaxies in the COSMOS field.
Although USmorph has demonstrated strong performance in applications involving large-scale datasets, its robustness under varying configurations has not yet been systematically quantified. For safe application to future deep- and wide-field surveys such as CSST, it is essential to ensure that the framework can deliver stable and scientifically reliable results. In this work, we conduct a comprehensive robustness analysis of USmorph. We evaluate its sensitivity to the architectural choices of the CAE, the effectiveness of the APCT in preserving morphological information under rotation, the optimization of the number of clusters $K$, and the stability of supervised classification under varying conditions. We also validate the scientific reliability of the final morphological classifications by examining the consistency between the predicted labels and galaxy evolutionary trends in physical parameter space. These parameters include the Sérsic index ($n$), effective radius ($r_{e}$), Gini coefficient ($G$), and the second-order moment of the brightest 20% of the flux ($M_{20}$). Through this analysis, we aim to prove the reliability of USmorph and provide a solid scientific basis for its application in future wide-field surveys.
[…]
Figure 1: The distribution in the $I_{\rm{mag}}$–redshift ($z$) parameter space of galaxies in the COSMOS field. Marginalized number distributions along the $I_{\rm{mag}}$ and $z$ dimensions are projected at the top margin and right margin, respectively. The subsample satisfying the selection criteria $0.2<z<1.2$ and $I_{\rm{mag}}<25$ is highlighted in red.
Figure 2: CAE architecture with symmetric encoder-decoder structure. Encoder: $100\times 100$ input undergoes two $5\times 5$ convolutional layers (16 channels throughout) interleaved with max pooling, producing $25\times 25\times 16$ features. The resulting feature maps are flattened to a 40-dimensional vector that encodes galaxy features.
II Observation and Sample Seletction
[…]
In addition to accurate photometric measurements, the COSMOS2020 catalog also provides estimates of various physical properties derived from spectral energy distribution (SED) fitting, such as photometric redshifts, stellar masses, and star formation rates. Two independent SED-fitting codes, EAZY (Bramme+2008) and LePhare (Ilbert+2006), were employed to estimate photometric redshifts. In this work, we adopt the redshifts derived from LePhare, as they demonstrate better reliability within our magnitude range of interest (see Figure 15 of Weaver+2022). The redshift estimation process utilizes a library of 33 galaxy templates constructed from the stellar population synthesis models of Bruzual+2003 and Ilbert+2009, along with a variety of dust attenuation prescriptions, including the starburst curve from Calzetti+2000, the SMC extinction curve (Prevot+1984), and two modified versions of the Calzetti+2000 law incorporating the 2175 Å bump. The final redshift ($z_{\text{LePh}}$) is defined as the median of the redshift likelihood distribution.
Figure 3: Denoising results using filters of different sizes. From left to right, the images show: the original noisy images; images denoised using Gaussian filtering ($\sigma=1$ and 22); images denoised with CAE filters of sizes $3\times 3$, $5\times 5$, and $7\times 7$; and images denoised using only a single convolutional layer. The denoising performance of the $5\times 5$ and $7\times 7$ filter sizes is quite similar; however, some small and faint structures in galaxies may be overlooked when using the $7\times 7$ filter, as indicated by the arrows. The white bar in the first panel indicates an angular scale of $1\arcsec$ ($\approx 33$ pixels).
We select the parent sample of galaxies from the COSMOS2020 catalog based on the following criteria:
  1. (1)
    $\rm lp_{type}=0$, to ensure that we are selecting galaxies rather than stars  [^1] For detailed parameter descriptions, please refer to: COSMOS 2020 “Farmer” Catalog.;
  2. (2)
    $\rm FLAG_{COMBINE}=0$, to ensure that the flux measurements are not affected by bright stars and that the target objects are not located at the edge of the image, thereby ensuring the reliability of redshift and stellar mass measurements;
  3. (3)
    $0.2<z<1.2$, to ensure that galaxy morphology is measured in the rest-frame optical wavelength range of approximately 3700–6800 Å.
  4. (4)
    $I_{\mathrm{mag}}<25$, to exclude galaxies that are too faint for reliable morphological measurements.
  5. (5)
    signal-to-noise ratio (S/N) greater than 5 and no abnormal pixels, to ensure the high-quality of source images.
99,806 galaxies were retained in the parent sample. Figure 1 shows the distribution of these galaxies in the $I_{\rm{mag}}$–redshift ($z$) parameter space.
III Analysis for The Unsupervised Method for Morphological Classification
This section introduces the workflow design and robustness validation of the UML module within the USmorph framework. The process of the UML module consists of three key stages: (1) using a CAE to reduce the dimensionality of raw galaxy images and extract key morphological features from high-dimensional image data; (2) utilizing APCT to resolve the rotational invariance problem in galaxy classification; (3) employing a bagging-based multi-clustering approach to cluster the galaxy images processed through the aforementioned steps.
III.1 data pre-processing
During the data pre-processing stage, we uniformly crop all images to $100\times 100$ pixels to retain sufficient structural information. Since the cropped size is related to the effective radii of galaxies in our sample, we use the GALAPAGOS software (Barden+2012; haubler+2022) to measure the effective radius of each galaxy. Subsequent statistical analysis shows that 98% of the sample galaxies have effective radii smaller than 50 pixels, confirming that the adopted $100\times 100$ pixel cropping size is sufficient for subsequent morphological classification tasks.
We further apply a max–min normalization pretreatment to each cutout. The flux of each pixel in float type is converted to nonnegative integers by the following procedure:
$f_{new}=\left\lfloor\frac{f-f_{min}}{f_{max}-f_{min}}\right\rfloor\times N,$ (1)
where $f_{max}$ and $f_{min}$ are the maximum and minimum fluxes of pixels, respectively. The floor function, $\lfloor x\rfloor$, gives the greatest integer less than or equal to $x$. NN is a given integer describing the degree of discretization, which is set as 500 in this paper. Our previous work (Zhou+2022) had shown that the choice of N does not have a significant influence on our results.
III.2 Convolutional Autoencoder Framework for Noise Reduction
Convolutional autoencoders (CAE; Masci+2011) represent an extension and refinement of traditional autoencoders (AE; Bourlard+1988). Unlike AEs that rely on fully connected layers for feature extraction, CAEs employ convolutional and deconvolutional layers that enable efficient learning of local structural features within images while preserving spatial information. Therefore, CAEs are widely used in image denoising, data compression, and semantic feature extraction (Krizhevsky+2012; Zhou+2022; Dai+2023). A CAE model consists of an encoder and a decoder. The encoder utilizes CNN to extract hierarchical features from the input data and progressively compresses the spatial information through a series of convolutional and max-pooling operations. The decoder reconstructs the denoised image from the compressed latent representation via transposed convolutional operations.
Figure 4: Denoising results obtained with different configurations. The first column shows raw images of five randomly selected galaxies; the second column shows the denoising results with the configuration used in this manuscript (Hidden dim=40); the third column shows the denoising results when the hidden dimension is reduced to 16; the fourth column shows the denoising results when 15 more training epochs are added; the fifth column shows the denoising results when using the binary cross-entropy loss function.
Figure 5: Left: Distribution of mean pixel-value differences for galaxy images before and after $90^{\circ}$ rotation, comparing cases with (blue) and without (grey) APCT. Each value represents the per-galaxy average of absolute differences $|\Delta I(x,y)|$ across all pixel positions. APCT results in near-zero differences, demonstrating strong rotational invariance. In contrast, non-APCT images exhibit significant deviations due to imperfect centrosymmetry. Right: Example galaxy processed with APCT under rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, showing that the APCT-processed images remain nearly unchanged, regardless of the rotation of the original input.
In this work, the input data dimension is set to $100\times 100$ (i.e., 10,000 pixels). The encoder comprises two layers, each containing a convolutional operation followed by a max-pooling step. The output is then flattened and compressed into a 40-dimensional vector via a fully connected layer in order to balance representational capacity and computational efficiency. The overall architecture of the proposed CAE is illustrated in Figure 2. The model is trained to minimize the reconstruction error between the input and the output images, quantified by a loss function, and to optimize the network parameters using the Adam optimizer (Kingma+2014). We employ the Rectified Linear Unit (ReLU; Agarap+2018) activation function and the Mean Squared Error (MSE; lehmann+2006) loss, defined as:
$\text{loss}=\frac{1}{10000n}\sum_{i=1}^{n}\sum_{j,k}^{100}(\hat{y}_{j,k}-y_{j,k})^{2},$ (2)
The loss is the mean squared error between all input and reconstructed pixels across all $n$ subsamples in a batch, where $y_{jk}$ and $\hat{y}_{jk}$ are the input and reconstructed pixel values at position $(j,k)$. In the training phase, a batch size of 8 is used, and the initial learning rate for the CAE is set to $3\times 10^{-4}$. The learning rate is adjusted dynamically using an exponential decay strategy (li+2019) to facilitate better optimization. The base number of training epochs is set to 32. If the validation loss plateaus, an additional 10 epochs are introduced to further enhance denoising performance, resulting in a total of 42 epochs.
We further conduct comparative experiments to validate the chosen CAE configuration. Figure 3 contrasts Gaussian (Gonzalez+2006) blurring with CAE-based denoising. The first column shows the raw images of five randomly selected galaxies; the second and third columns present Gaussian filtering results with standard deviations of $\sigma=2$ and $\sigma=3$, respectively; the fourth to sixth columns display denoising outcomes using CAEs with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$. As shown in Figure 3, Gaussian blurring tends to smear structural features of the galaxies, whereas the CAE approach preserves these features more effectively. For instance, Gaussian filtering significantly increases the apparent size of the first galaxy, while certain features in the fourth and fifth galaxies are excessively smoothed. Kernel size considerably influences denoising performance: smaller kernels (e.g., $3\times 3$) retain more residual noise, while larger kernels (e.g., $7\times 7$) suppress faint structures, as indicated by the red arrows. The seventh column presents results from a CAE using a single $5\times 5$ convolutional layer, which performs comparably to a two-layer CAE with $3\times 3$ kernels. This suggests that increasing network depth can be functionally similar to enlarging the receptive field via larger kernels. However, we settle on an encoder comprising two convolutional layers with $5\times 5$ kernels since very deep networks may also overlook faint features.
When selecting the output dimensionality of the CAE, a trade-off must be made between computational efficiency and the preservation of image features: excessively high dimensionality increases the computational cost of downstream clustering tasks, while overly low dimensionality may result in the loss of key morphological features of galaxies. Zhou+2022 found that increasing the dimensionality to 100 preserves all features but only provides limited improvement in classification performance. Therefore, we adopt a 40-dimensional representation in this work, which maximizes computational efficiency while retaining the essential morphological information of galaxy images. In Figure 4, the first column shows the raw images of 5 randomly selected galaxies; the second column presents the denoised results obtained with the adopted CAE configuration (hidden dim=40); and the third column shows the reconstruction when the latent dimensionality is set to 16. A comparison clearly demonstrates that an excessively small dimensionality leads to substantial feature loss.
We further conduct several other comparative experiments: the fourth column displays the denoised results after training for an additional 25 epochs beyond loss stabilization (i.e., a total of 15 extra epochs), which show no evident improvement over the second column; the fifth column shows the results obtained when an alternative loss function is used (binary cross-entropy), which fails to visually outperform the adopted configuration because it loses certain galaxy details. Figures 3 and 4 further validate the robustness of our parameter choices.
III.3 The APCT Preprocessing Strategies
Convolutional Neural Networks (CNNs) possess translation invariance but lack rotational invariance in image classification. This deficiency can cause misclassifications of rotated galaxy images that share identical morphology (yao+2019). Traditional approaches, such as “rotated image augmentation,” enhance rotational robustness by substantially expanding the training set (e.g., adding 2288 rotated versions per image; Dominguez+2018). However, this method lacks efficiency and relies on similarity between the angular distributions of the training and test sets. It fails to handle the scenario where galaxy orientations are randomly distributed in real observations.
To ensure rotationally invariant classification of galaxy morphology, we employ Adaptive Polar Coordinate Transformation (APCT; Fang+2023). Conventional polar coordinate transformations utilize a fixed polar axis, which can induce horizontal shifts in the transformed images when rotating the original images. In contrast, APCT defines the coordinate system using a rotationally invariant polar axis instead of a fixed one. The technique first identifies the brightest and darkest pixels within the image to determine the direction of the polar axis. The axis is then rotated counterclockwise in $0.05$ rad steps. Subsequently, all pixels are mapped (“stacked”) onto the polar coordinate space along the rotated axis. Finally, a mirroring operation is applied to the image to accentuate central features. This process enhances the model’s rotational invariance and simultaneously improves the saliency of morphological features within the image.
Figure 5 illustrates the distribution of the mean pixel-value differences before and after a 90 rotation of galaxy images, with and without APCT. The results show that, with APCT, the majority of images exhibit minimal pixel-value differences, whereas without APCT, the differences are significantly larger due to the lack of perfect central symmetry in galaxies. This result demonstrates that the APCT effectively eliminates the effect of rotation on pixel distribution while preserving the core features of the image by converting Cartesian coordinates to polar coordinates. By transforming the problem of rotation invariance into one of translation invariance, the APCT can significantly enhance the rotational classification robustness of CNNs, which will be further validated in  V.1 via more tests.
III.4 Voting strategy and Post-hoc label alignment
Hybrid clustering voting boosts (breiman+1996) galaxy classification robustness and accuracy by integrating multiple models. Given that galaxies frequently exhibit transitional morphologies during their evolution, and different algorithms employ distinct similarity criteria, individual methods struggle with ambiguous samples from evolutionary sequences. Our Bagging-based multi-model voting approach minimizes errors by excluding samples with significant inter-model disagreement (Zhou+2022).
The specific implementation combines three clustering algorithms: K-means Clustering (K-means; Hartigan+1979), Hierarchical Balanced Iterative Reduction and Clustering Algorithm (BIRCH; Zhang+1996), and the Agglomerative Clustering algorithm (Agg; Murtagh+1983; Murtagh+2014). This collective approach reduces the inherent errors of individual models. Samples achieving unanimous agreement among all three algorithms constitute the final consensus set, while discordant samples are excluded. The schematic diagram is shown in Figure 4 of Zhou+2022.
We choose these three clustering algorithms due to their complementary properties: K-means is particularly effective for convex datasets; Agg captures hierarchical structures and performs well on concave distributions; and Birch efficiently handles noisy and large-scale data. Specifically, K-means rapidly partitions data into initial clusters based on distance metrics; Agg identifies hierarchical relationships within the data; and Birch processes large datasets efficiently by constructing a clustering feature tree.
By integrating the clustering results of all three algorithms through a voting mechanism, we effectively minimize the bias introduced by individual algorithms in the galaxy morphology classification task, substantially improving clustering purity (Zhou+2022; Dai+2023; liu+2023). In contrast, other algorithms, such as density-based spatial clustering of applications with noise (DBSCAN;Deng+2020) and Mean shift clustering (Mean-shift; Comaniciu+2002), present challenges for application in astronomical survey projects with massive data due to their parameter sensitivity, cumbersome parameter adjustment, and insufficient efficiency in processing large-scale data.
After determining the voting strategy, we adopt a post-hoc label alignment approach (liu+2023), which consists of two steps: over-clustering and visual category mapping. In unsupervised image classification tasks, clustering partitions samples solely based on feature similarity, and its inherent ”semantic deficiency” makes the results difficult to directly correspond to physical research objectives. Therefore, the reasonable selection of the number of clusters is crucial. A common practice is to preset a relatively large number of categories based on scientific goals, allowing the unsupervised algorithm to fully capture fine-grained feature differences, and then assign semantic information to the clustering results through a small number of visual annotations.
In this study, we set $k=50$ based on the following considerations. Firstly, $k=50$ is sufficient to ensure that each subcategory is highly homogeneous in morphology. Our tests demonstrate that a set of 50 clustering samples can typically be annotated within five minutes, without imposing excessive manual effort. Secondly, a larger number of categories helps to characterize subcategory structures and identify special celestial bodies (such as tidal tails, strong gravitational lenses, and littlered dots), thereby providing deeper support for subsequent scientific analysis [^2] See Figure 4 of Song+2024 for Post-hoc label alignment..
The 63,202 galaxies receive inconsistent votes with different clustering algorithms, accounting for 64% of the samples. This relatively high fraction stems primarily from discrepancies in the voting results among different clustering algorithms. This indicates that these galaxies exhibit ambiguous or transitional morphological features, which preclude consistent classification. After excluding samples with inconsistent votes, a total of 36,604 galaxies were successfully classified into 50 distinct groups. Ultimately, the remaining 36,604 galaxies were categorized into five morphological classes: spherical (SPH), early-type disk (ETD), late-type disk (LTD), irregular (IRR), and unclassified (UNC).
Figure 6 presents the results of UML clustering along with some representative samples. The effectiveness of these 50 clusters will be validated in subsequent experiments. The final sample comprises: 8,233 SPH, 5,322 ETD, 6,320 LTD, 9,468 IRR, and 7,261 UNC galaxies.
Figure 6: An overview of the UML dataset. Left:: cutouts of galaxies selected from different categories, i.e., SPH, ETD, LTD, IRR, and UNC galaxies, are shown from left to right. Right: the UML dataset illustrates the statistical distribution features of 50 galaxy groups (panel (a)) and the corresponding five-class merger label distributions (panel (b)).
Figure 7: The distribution ratio of different types of galaxies in the training and validation sets remains consistent. This consistency ensures the reliability and generalization ability of the model evaluation results.
Figure 8: The distribution of training (red) and validation (blue) set data across various physical parameters. The consistent distribution between the training and validation sets illustrates the reliability of our methodology.
IV Analysis for The SML Models and the Experiment Settings
To achieve automated galaxy morphology classification, we utilize a training set of 36,604 reliably labeled sources from UML clustering. Using this dataset, we train and optimize the SML algorithm. This section details the implementation specifics and experimental configuration of the SML algorithm.
IV.1 The GoogLeNet Model
The GoogLeNet Model (szegedy+2015), a deep convolutional neural network, employs the innovative Inception module, which features a multi-scale convolutional parallel architecture. This design elegantly achieves a balance between network depth, width, and computational efficiency, demonstrating exceptional performance across a wide range of application scenarios. The GoogLeNet architecture uses nine sequential inception modules, each applying parallel $1\times 1$, $3\times 3$, and $5\times 5$ convolutions. This design captures broad image areas while preserving fine details. With increasing network depth and width, GoogLeNet can achieve greater parameter efficiency, a reduction in total parameters, and improvements in both performance and computational efficiency. In a comparative analysis, Fang+2023 demonstrated that GoogLeNet exhibited the highest stability among three deep learning frameworks for galaxy image classification.
IV.2 Experiment Setting
We utilize 36,604 reliably labeled samples obtained from UML clustering as the training set to classify the remaining 63,202 galaxies using SML. In principle, the entire UML dataset should be employed for training to enable the algorithm to fully exploit all available information, but this may lead to overfitting. Taking the large sample size, we divide the UML dataset into training (Tra) and validation (Val) sets at a 9:1 ratio.
We further verify that key feature distributions were balanced between the two sets to ensure the model’s generalization capability and the reliability of performance evaluation, as well as to avoid potential biases caused by discrepancies in data distribution. As illustrated in Figure 7, the distributions of galaxy morphological types are consistent between the training and validation sets. Similarly, Figure 8 shows that their distributions in the parameter space of physical properties—such as stellar mass ($M_{\ast}$), redshift ($z$), and other morphological parameters are highly similar (see Section V.3). The Kolmogorov–Smirnov (K–S; Kolmogorov+1933) test results (Table 1) show that all the examined parameters yield $p$-values greater than 0.05, indicating that the training and validation sets follow statistically consistent distributions.
Table 1: Results of Partial K-S Test
parameter Parameters
stellar mass redshift $\rm R_{e}$ n
p-value 0.82 0.08 0.34 0.64
NUV-r r-J $\rm I_{mag}$ Gini
p-value 0.34 0.33 0.73 0.64
  • With $p>0.05$, indicating that the training and validation sets follow the same distribution for the corresponding parameter.
Table 2: The confusion matrix
Galaxy Type Recall Precision F1-Score
SPH 98.8% 99.4% 99.0%
ETD 95.4% 93.8% 94.6%
LTD 91.6% 91.0% 91.3%
IRR 93.5% 91.0% 92.2%
UNC 91.8% 96.2% 94.0%
  • Note: The overall accuracy of the GoogLeNet model is above 94%, which proves the robustness of the galaxy classification task.
Figure 9: Panels (a), (b), and (c) show the t-SNE diagrams of galaxies in the samples obtained using UML, SML, and the combined method, respectively. Panels (d), (e), and (f) present the corresponding class-wise contours for each of the three diagrams, enclosing 20%, 50%, and 70% of the respective samples. The clear boundaries between different galaxy types in the t-SNE maps indicate that the classification algorithms (UML, SML, and the combined model) exhibit strong robustness.
Figure 10: Panels (a) and (b) show the distribution of galaxies in rest-frame NUVrNUV-r vs. rJr-J color space and in the $G-M_{20}$ parameter space, respectively. The contours enclose 70% of the galaxies in each subclass. The scattered points represent 500 galaxies randomly selected from each subclass. Colored diamonds indicate the median values of the distributions, and the associated error bars represent the standard deviations. Panels (c) and (d) display the distributions of Sérsic index ($n$) and effective radius ($r_{\text{e}}$), respectively. In the box plots, the boxes span the interquartile range (IQR) from the first to the third quartile, the central black line marks the median, and the whiskers extend to the minimum and maximum values.
We take a batch size = 32 during the supervised training phase. The learning rate is reduced to 0.0001, which facilitates stable convergence of the model. To ensure stable and efficient convergence during training, the GoogLeNet model is optimized using stochastic gradient descent (SGD; Sutskever+2013) with momentum. We use a maximum of 300 training epochs. If the overall accuracy does not show a significant improvement after 150 epochs, the training will be terminated to mitigate overfitting and avoid waste of computational resources. We perform training on a server with an NVIDIA A10 GPU. Our configuration achieves a balance between training efficiency and generalization capability within the constraints of available video memory capacity. The overall parameter setup emphasizes the integration of performance efficacy and system stability.
IV.3 Performances of SML Models
To verify the effectiveness of the GoogLeNet model in galaxy image classification, we employ Precision, Recall, F1-Score, and Accuracy as performance evaluation metrics. All the aforementioned metrics are calculated based on the counts of true positives (TP), false positives (FP), true negatives (TN), and false negatives (FN), with the formulas given below:
$\textit{Precision}=\frac{TP}{TP+FP}$ (3)
$\textit{Recall}=\frac{TP}{TP+FN}$ (4)
$\textit{F1-Score}=2\times\frac{\textit{Precision}\times\textit{Recall}}{\textit{Precision}+\textit{}{Recall}}$ (5)
$\textit{Accuracy}=\frac{TP+TN}{TP+FP+TN+FN}$ (6)
Table 2 shows GoogLeNet’s excellent performance, achieving an overall accuracy of 94%. To evaluate the model’s generalization ability and detect potential overfitting, we allocate half of the validation set as an independent test set. Simultaneously, we assess the stability of model evaluation across different data partitions by repeated random subsampling. The accuracy (Acc) on both the validation and test sets is largely unaffected by variations in the split ratio, as shown in Table 3. Our tests show that the GoogLeNet model consistently produces stable feature representations. These results collectively provide evidence that no overfitting was observed during training.
Table 3: Model Accuracy Across Different Data Split Ratios
Division (Tra Set) TA (Val Set) TA (Test Set) TA
$7:1.5:1.5$ 100% 93.8% 93.4%
$8:1:1$ 100% 93.9% 93.8%
$9:0.5:0.5$ 100% 95.3% 94.1%
  • Note: To obtain an unbiased evaluation, we introduce a test set to ultimately verify the model’s true generalization ability and detect overfitting.
V RESULTS AND DISCUSSION
In this section, we demonstrate the effectiveness of the unsupervised processing strategy, present the complete results of galaxy morphological classification, and verify their effectiveness through a series of tests.
V.1 The effectiveness of CAE and APCT
To quantitatively evaluate the effectiveness of data processing, we performed Bagging clustering analysis on two types of images: original images and processed images. Based on the reliable labels obtained from the clustering analysis, we computed the overall classification accuracy of the GoogLeNet model on the validation set under test conditions including image rotation angles of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$. The results are summarized in Table 4.
Table 4 demonstrates that data processing effectively enhances the classification performance of GoogLeNet. When the validation images are not rotated, the performance gap between models trained with and without data processing is relatively small. This is because the angular distribution difference between the validation set (randomly sampled from the UML dataset) and the training set is minimal. In contrast, rotating the validation images introduces a distribution shift in orientation between the training and validation sets. Under such conditions, the accuracy of the model trained without data processing decreases significantly, while the model trained on processed images maintains stable performance. These results align with Fang+2023 and Zhang+2024, further validating the effectiveness of the adopted data processing strategy.
Table 4: The validation accuracy under different data processing methods and at various rotation angles.
Method dataset Count 90° 180° 270°
Raw image 39,896 90.5% 86.9% 88.9% 86.5%
Processed image 36,604 94.4% 93.9% 93.7% 93.8%
Figure 11: The precision of each category of galaxies in the parameter space of the validation set (red: SPH, blue: ETD, yellow: LTD, green: IRR, and grey: UNC). Panels (a), (b), (c), (d), (e), and (f) represent the precision of galaxies on parameters such as Sersic index ($n$), radius ($r_{e}$), Gini($G$), M20, and NUVrNUV-r and rJr-J colors. We define the parameter range to encompass all galaxy types and calculate the median precision of each parameter interval within a metric for its classification precision. The parameter range includes all galaxy types, and the median precision within each parameter interval is calculated as a metric for its classification precision. The label gives the median precision of this type of galaxy within the entire parameter range. The precision of the model for classifying various categories of galaxies is independent of galaxy morphology and physical parameters. We find no significant difference or trend in the precision of each category in the parameter space, indicating that the False Positives do not rely on galaxy morphological and physical parameters.
Figure 12: Same as Figure 11 but for the recall metric. Similar to the results of precision, we find no significant difference or trend in the recall of each category in the parameter space, indicating the False Negatives do not rely on galaxy morphological and physical parameters. The recall of each category of galaxies in the parameter space of the validation set. The sequence and symbols are the same as Figure 11. The model’s recall across different galaxy categories is independent of morphology and physical parameters, indicating that False Negatives show no significant dependence on these factors.
V.2 Overall morphological classification results and t-SNE Visualization
By combining UML and SML, we successfully perform morphological classification of 99,806 galaxies in the COSMOS field, as shown in Table 5. The final classification results comprise 15,326 SPH, 17,445 ETD, 20,931 LTD, 29,123 IRR, and 16,981 UNC. We then visualize the clustering results with the t-distributed stochastic neighbor embedding (t-SNE; van+2008), which projects high-dimensional data into a lower-dimensional space (e.g., 2D or 3D). As shown in Figure 9, t-SNE effectively maps the high-dimensional galaxy samples into a 2D plane while preserving both local similarities and global structural features. t-SNE visualization allows us to qualitatively evaluate the classification performance of UML, SML, and the combined model. Figure 9 demonstrates that all three methods produce well-separated class boundaries in the feature subspace: galaxies with similar features are accurately mapped to neighboring positions in the low-dimensional space, whereas galaxies with significant feature differences exhibit clear distributional separation. These results confirm the effectiveness of our classification approach.
[…]
Table 5: The number of galaxies classified into different types
TYPE SPH ETD LTD IRR UNC TOTAL
UML 8233 5322 6320 9468 7261 36,604
SML 7093 12,123 14,611 19,655 9720 63,202
TOTAL 15,326 17,445 20,931 29,123 16,981 99,806
V.3 Comparisons of Galaxy Properties
To verify the consistency between the classification results and the known structural patterns of galaxies, we analyze the classification results using the physical parameters of galaxies in Figure 10. To ensure the accuracy and reliability of the parameter measurements, the analysis is restricted to massive galaxies ($M_{*}>10^{9}~M_{\sun}$). Galaxies classified as UNC are excluded because their morphological parameters are difficult to measure due to the relatively low signal-to-noise ratios (S/N). In Figure 10, the SPH, ETD, LTD, and IRR categories are colored in red, blue, yellow, and green, respectively.
The panel (a) panel shows the distribution of galaxies in the rest-frame NUVrNUV-r versus rJr-J color space. The NUVJNUV-J diagram is an effective diagnostic tool for distinguishing star-forming from quiescent galaxies; the wedge-shaped region is dominated by quiescent galaxies, while galaxies outside this region are considered star-forming (Williams+2009). Typically, ETD, LTD, and IRR galaxies primarily locate in the star-forming region, whereas most SPH galaxies are identified as quiescent.
The panel (b) panel presents the Gini coefficient ($G$) against the $M_{20}$ parameter (Lotz+2004; Lotz+2008). The Gini coefficient quantifies the non-uniformity of the surface brightness distribution, with higher values indicating stronger non-uniformity; $M_{20}$ is the second moment of the brightest 20% of the galaxy’s light, with more negative values implying a higher degree of central concentration.
The bottom panels show raincloud plots for the distributions of the Sérsic index $n$ (Panel (c)) and the effective radius $r_{e}$ (Panel (d)). Their definitions of these parameters are given in previous work (Zhou+2022; Song+2024).
Figure 10 shows that our classification results are consistent with the expected evolutionary trend: along the sequence, from IRR to SPH types, (a) galaxies migrate from the star-forming region to the quiescent region, (b) their light distribution becomes more centrally concentrated, (c) galaxies gradually become more bulge-dominated, and (d) their morphology evolves from extended to more compact structures.
We use the metrics of precision and recall to quantify the classification robustness of the GoogLeNet model across different parameter spaces and to investigate the relationship between FP and FN. Precision measures the accuracy of positive predictions, which increases as false positives decrease. Recall assesses the coverage of true positives prior to detection, which increases as false negatives decrease. Figures 11 and 12 show the classification precision and recall, respectively, for the five galaxy classes (SPH, ETD, LTD, IRR, UNC) on the validation set. The results show no significant correlation between these metrics and the morphological parameters of the galaxies. Both precision and recall remain above 90% for all classes. Figures 11 and 12 demonstrate that the distributions of true positives and false positives are independent of galactic characteristics. This confirms that the model delivers highly accurate and stable classification outputs across all categories, effectively avoiding biases introduced by specific sample properties.
VI Summary and outlooks
This paper presents a systematic analysis of the robustness of the hybrid machine learning framework USmorph, which integrates UML and SML for the automated classification of large-scale galaxy morphologies. The framework is validated with a sample of nearly 100,000 I-band galaxies ($0.2<z<1.2$, $I_{\mathrm{mag}}<25$) drawn from the COSMOS field.
In the unsupervised learning phase, we verify that: (1) A CAE with two network layers, $5\times 5$ convolutional kernels, and a 40-dimensional latent representation effectively suppresses background noise while maximally preserving key galactic structures. (2) The APCT is proven indispensable for achieving rotational invariance. APCT transforms rotational invariance into translational invariance, thereby conserving computational resources and enhancing the efficiency of downstream tasks. (3) Bagging multi-clustering strategy yields high-purity consensus label classes. Setting the number of clusters to K=50 balances classification granularity with annotation efficiency and facilitates subsequent exploration of special objects (e.g., tidal tails, gravitational lenses, and little red dots) through subclass refinement.
In the supervised learning phase, the GoogLeNet model achieves a consistent overall classification accuracy exceeding 94% across different dataset splits, demonstrating remarkable stability. Further validation involves t-SNE visualization and physical morphological parameters in the classification study. The results indicate that morphology classification based on USmorph exhibits distinct low-dimensional clustering boundaries in t-SNE visualization and maintains a high degree of consistency with the evolution of galaxy morphological parameter features. Additionally, both true positive and false positive samples displayed unbiased distributions within the parameter space, further substantiating the high reliability and stability of this classification method.
In future work, we will further enhance the automatic galaxy morphology classification framework by developing more effective large-scale model coding and dimensionality reduction techniques. We will also assess the stability and effectiveness of these algorithms (e.g., Ying et al. 2025, in preparation). We aim to improve the accuracy and generalization performance of deep learning models. This framework is anticipated to be applied to upcoming multi-band deep-field surveys, such as the imaging data expected from CSST, thereby advancing the study of galaxy morphology.
Data Availability
Our code, dataset, and trained model weights can be found at https://github.com/IAAA-246011/USmorph.
This work is supported by the National Natural Science Foundation of China (NSFC) (grant Nos. 12233008 and 12573012, and Project No. 12503011), the National Key R&D Program of China (grant No. 2023YFA1608100), the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550200), the Cyrus Chun Ying Tang Foundations, and the 111 Project for “Observational and Theoretical Research on Dark Matter and Dark Energy” (B23042), and the China Manned Space Program with grant No. CMS-CSST-2025-A04. S.L. acknowledges the support from the Key Laboratory of Modern Astronomy and Astrophysics (Nanjing University) by the Ministry of Education. The numerical calculations in this paper have been done on the platform of the High Performance Computing of Anqing Normal University.